In a 9 inch square cake pan and a 10 inch square cake pan, what is the difference in volumn each pan will hold? assume each pan is 3 inches high. step by step, please.
Answer by rapaljer(4671)

V = LWH

V1=10*10*3 = 300 cubic inches
V2= 9*9*3= 243 cubic inches

The difference = 300-243 = 57 cubic inches.

Dr. Rapalje
Seminole State College of Florida

P.S. See my website by clicking on my tutor name "rapaljer". The link there will take you to my homepage. About halfway down my homepage, look for "Topics from Geometry," and look for the link "Perimeters, Areas, and Volumes."